my problem was 4(2x-3)+8(x-4)=2x+6 and the answer that i got was 5.6 how would that be in a fraction?
Answer by orca(409)

The solution I got:
4(2x-3)+8(x-4)=2x+6
8x - 12 + 8x - 32 = 2x + 6
16x - 44 = 2x + 6
14x = 50
x = 50/14 = 25/7
Or written in decimal x = 3.57.